x^2-9÷x-3=-3 at x=3 but for me this =0÷0 which is N.D.

Grade:12

1 Answers

Vikas TU
7 years ago
Yes it is 0/0 indeterminant form.
And thus we need to solve this by factorisation if it is reducing further .
Therefore,
(x+3) (x-3) / (x -3) = > x+3
at x = 3  the value it gives is  = 3 + 3 = 6.